Prove geometrically, (a-b)(a+b) =(a^2)-(b^2)
Answer by ptth176(50)

(a-b)(a+b) =(a^2)-(b^2)
(1) (2)
(1)=>a^2-ab+ab-b^2=a^2-b^2
=>(1)=(2)